(a)               A chemical preparation has a theoretical yield of 16.50g, the actual yield was 12.375g.  Express the actual yield as a percentage of the theoretical yield.


(b)               The experimental value of the water of hydration of copper sulfate was calculated as 36.1%.  What mass of a sample of 18g of hydrated copper sulphate is water?  Write your answer correct to three significant figures.



(c)       A reaction produces a compound at a rate of 12 gmin-1.  A catalyst raises this rate to 16.5 g min-1. What is the percentage increase?


1
Expert's answer
2021-03-07T17:07:25-0500

(a) "\frac{12.375}{16.50}=0.75=75\%."


(b) "18*0.361=6.498\approx6.50\; g."


(c) "\frac{16.5-12}{12}=0.375=37.5\%."